find two consecutive even integers such that the larger,added to eight times the smaller, equals 110.i have tried to work out this problem but its very confusing for me.This is how i tried it x+2+8=110,x+10=110.then i minused the 10 from the 110.And my answer was x=100.

Answer by checkley77(12569)

x+2+8=110 YOU FORGOT THAT THE Y MULTIPLIER FOR THE 8.
LET X & X=2 BE THE 2 INTEGERS.
(X+2)+8X=110
9X+2=110
9X=110-2
9X=108
X=108/9
X=12 ANS.
PROOF:
12+2=14 ANS. FOR THE LARGER INTEGER.
14+8*12=110
14+96=110
110=110